given the function of using C = 700 = 0.75yd, T = 100, I = 200, R = 50, G = 300 X = 250 and M = 200 + 0.1y. Calculate the equilibrium income
1
Expert's answer
2019-10-14T08:55:29-0400

The equilibrium income is:

Y = C + I + G + NX = 700 + 0.75Yd + 200 + 300 + 250 - (200 + 0.1Y) = 700 + 0.75(Y - 100) + 750 - 200 - 0.1Y = 1175 + 0.65Y,

0.35Y = 1175,

Y = 1175/0.35 = 3357.14.